OpenStudy (anonymous):

what is the integration of 1/1-2.v^2 dv ??

OpenStudy (irishboy123):

\(\int {1 \over 1-2v^2} dv \) ?? if so \( {1 \over 1-2v^2} = {1 \over (1-\sqrt{2}v)(1+\sqrt{2}v)}\) then partial fractions and log integrals....